# Split Two-Periodic Aztec Diamond

### v4

@@ 2.2. Comparison to the Two-Periodic Aztec Diamond
Before we analyze the behavior of the model, we first want to comment on how the form of the correlation kernel relates to the correlation kernel of the typical two-periodic Aztec diamond. From the definition of the model, we can recover the two-periodic Aztec diamond (with two-periodic parameter $\varepsilon$) from the split two-periodic Aztec diamond by setting $\alpha=\beta=\varepsilon$. The correlation kernel of the two-periodic Aztec diamond, as formulated in [3], can be stated using the notation presented in this paper as
Theorem ([3], Theorem 5.2).
Let $-N\leq\xi,\xi^{\prime}\leq-1$ and $0<m<N$. The two-periodic Aztec diamond of size $2N$ has correlation kernel given by
(17) $\Big{[}\mathbb{K}_{TP,N}^{\varepsilon}(4m^{\prime},2\xi^{\prime}+j;4m,2\xi+i)\Big{]}_{i,j=0}^{1}=-\frac{\mathbb{I}_{m>m^{\prime}}}{2\pi\emph{i}}\oint_{\gamma_{0,1}}\frac{\text{d}z}{z}z^{\xi^{\prime}-\xi}\Phi_{\varepsilon}(z)^{m-m^{\prime}}\\ +\frac{1}{(2\pi\emph{i})^{2}}\sum_{k=1,2}\oint_{\gamma_{1}}\text{d}w\oint_{\gamma_{0,1}}\frac{\text{d}z}{z(z-w)}\frac{w^{\xi^{\prime}+N}(z-1)^{N}}{z^{\xi+N}(w-1)^{N}}r_{\varepsilon,1}(w)^{\frac{N}{2}-m^{\prime}}r_{\varepsilon,k}(z)^{m-\frac{N}{2}}F_{\varepsilon,1}(w)F_{\varepsilon,k}(z)$
@@
(18) $\Big{[}\mathbb{K}_{N}(4m^{\prime},2\xi^{\prime}+j;4m,2\xi+i)\Big{]}_{i,j=0}^{1}=\Big{[}\mathbb{K}_{TP,N}^{\varepsilon}(4m^{\prime},2\xi^{\prime}+j;4m,2\xi+i)\Big{]}_{i,j=0}^{1}\\ +\frac{1}{(2\pi\textrm{i})^{2}}\sum_{k=1,2}\oint_{\gamma_{1}}\text{d}w\oint_{\gamma_{0,1}}\frac{\text{d}z}{z(z-w)}\frac{w^{\xi^{\prime}+N}(z-1)^{N}}{z^{\xi+N}(w-1)^{N}}r_{\varepsilon,l}(w)^{\frac{N}{2}-m^{\prime}}r_{\varepsilon,k}(z)^{m-\frac{N}{2}}\\ \times\frac{2}{1+2g_{\alpha,\beta}(w)}F_{\varepsilon,l}(w)F_{\bar{\varepsilon},\bar{l}}(w)F_{\varepsilon,\bar{l}}(w)F_{\varepsilon,k}(z)\$
@@
$(l,\bar{l})=\begin{cases}(2,1)&\text{if }m,\,m^{\prime}<\nicefrac{{N}}{{2}}\\ (1,2)&\text{if }m,\,m^{\prime}>\nicefrac{{N}}{{2}}\end{cases}$
@@ 2.3. Classification of the Macroscopic Regions
Before we compute the explicit local asymptotics of the model, we would like to introduce the saddle functions associated with the kernel of the model. In doing so, we will consider the behavior of the model around some asymptotic coordinate. Let $(x,y)\in(0,\nicefrac{{1}}{{2}})\cup(\nicefrac{{1}}{{2}},1)\times(-1,0)$, where we omit the case where the asymptotic coordinate is exactly on the interface of the model. By letting $(m,\xi)=(xN,yN)$, the vertex $(4m,2\xi+i)$ has asymptotic coordinate $(x,y)$. Now we define the functions,
@@
The saddle functions $\psi_{\varepsilon,k}(z)$ are, up to a scalar and change of parameters, the saddle functions defined and analyzed in [11].
In other words, most of the analytic work we need in order to prepare our integrals for saddle point methods was accomplished in [11]. We state the saddle functions in [11, p. 15] for comparison.
@@
$\lambda_{1,2}(z)=\frac{\left((\alpha+\beta)z\pm\sqrt{z(z+\alpha^{2})(z+\beta^{2})}\right)^{2}}{z(z-1)^{2}}$
and $-1<\xi_{1},\xi_{2}<1$ are parameters that define the asymptotic coordinate, albeit different from our $x$ and $y$. Note that the above $\alpha$ and $\beta$ are from the convention in [11] and not the same $\alpha$ and $\beta$ presented in this text. Letting $\alpha=\varepsilon$ and $\beta=\varepsilon^{-1}$ it is easy to check that,
@@
As long as $(x,y)$ is not on some boundary between regions, the function $\Psi(z)$ has a total of four simple saddle points. The location of these saddle points on the Riemann surface classify the macroscopic region of the model.
@@
    If there are two saddles points on either copy of the positive real axis, then $(x,y)\in\mathfrak{F}$, the frozen region of the model.
@@
    If there are two complex saddle points (on either sheet of the Riemann surface), then $(x,y)\in\mathfrak{R}$, the rough region of the model.
@@
    If all four saddle points are located on either copy of the negative real axis away from the branch cuts, then $(x,y)\in\mathfrak{S}$, the smooth region of the model.
This is the same definition of regions presented for the typical two-periodic model in [11, Definition 2.9]. The saddle function, $\Psi(z)$, can also be used to determine the boundaries between the macroscopic regions. Any point $(x,y)$ where saddle points coalesce is on some boundary between the macroscopic regions. Additionally, due to the piece wise nature of the saddle functions, we must check for transitions along the line $x=\nicefrac{{1}}{{2}}$. Figure 7 depicts an example of the macroscopic boundaries for a split two-periodic model.
@@
For either half of the model, the definition and boundaries between macroscopic regions agree with the typically two-periodic case. [^4] See [11] for a similar classification in the two periodic case. In the section below we will show how this follows from the local asymptotics of the correlation kernel.
@@ 2.4. Asymptotic Behavior
Our goal in this section is to summarize the local asymptotics of the model, so we will assume the two coordinates $(4m^{\prime},2\xi^{\prime}+j)$ and $(4m,2\xi+i)$ are on the same side of the interface and that they are asymptotically close. With these considerations in mind, we can start from the phrasing of the kernel given in equation (18).
This work will only discuss the asymptotic results of the model away from the interface. To begin, we define the coordinates in the following manner
@@
where we assume $x\neq\nicefrac{{1}}{{2}}$, and $x_{1}$ and $x_{2}$ are small enough in magnitude so that the coordinates $(4m,2\xi+i)$ and $(4m^{\prime},2\xi^{\prime}+j)$ lie on the same side of the interface. In general, we will assume that $x_{1},\,x_{2},\,y_{1},$ and $y_{2}$ are $o(N)$. Under these considerations, we rewrite correlation kernel in terms of the saddle functions,
(23) $\Big{[}\mathbb{K}_{N}(4m^{\prime},2\xi^{\prime}+j;4m,2\xi+i)\Big{]}_{i,j=0}^{1}=\Big{[}\mathbb{K}_{TP,N}^{\varepsilon}(4m^{\prime},2\xi^{\prime}+j;4m,2\xi+i)\Big{]}_{i,j=0}^{1}\\ +\frac{1}{(2\pi\textrm{i})^{2}}\sum_{k=1,2}\oint_{\gamma_{1}}\text{d}w\oint_{\gamma_{0,1}}\frac{\text{d}z}{z(z-w)}\frac{2}{1+2g_{\alpha,\beta}(w)}F_{\varepsilon,l}(w)F_{\bar{\varepsilon},\bar{l}}(w)F_{\varepsilon,\bar{l}}(w)F_{\varepsilon,k}(z)\\ \times\exp\left[N\left(\psi_{l}(w;x,y)-\psi_{k}(z;x,y)\right)+\varphi_{\varepsilon,l}(w;x_{2},y_{2})-\varphi_{\varepsilon,k}(z;x_{1},y_{1})\right]$
@@
For brevity we will let
(25) $\Big{[}I^{\varepsilon}_{l,k}(4m^{\prime},2\xi^{\prime}+j;4m,2\xi+i)\Big{]}_{i,j=0}^{1}\\ =\frac{1}{(2\pi\textrm{i})^{2}}\oint_{\gamma_{1}}\text{d}w\oint_{\gamma_{0,1}}\frac{\text{d}z}{z(z-w)}\frac{2}{1+2g_{\alpha,\beta}(w)}F_{\varepsilon,l}(w)F_{\bar{\varepsilon},\bar{l}}(w)F_{\varepsilon,\bar{l}}(w)F_{\varepsilon,k}(z)\\ \times\exp\left[N\left(\psi_{l}(w;x,y)-\psi_{k}(z;x,y)\right)+\varphi_{\varepsilon,l}(w;x_{2},y_{2})-\varphi_{\varepsilon,k}(z;x_{1},y_{1})\right]$
for $k=1,2$. So we may write the kernel, briefly, as
@@
Ultimately, our goal is to show that the contributions from the contour integrals $I^{\alpha}_{2,1}$ and $I^{\alpha}_{2,2}$ tend to 00 as $N\to\infty$. This will justify the definition of the macroscopic boundary given in Definition 2.1. The order of decay depends on the region of the model of the asymptotic coordinate. We will find that, in certain regions, these additional terms affect the second order term of the local asymptotics of $\mathbb{K}_{N}$.
@@
Assuming the coordinate definition given in equations (21) and (22), this means that the local asymptotic behavior of the contour integral $I^{\varepsilon}_{2,2}$ when $x<\nicefrac{{1}}{{2}}$ is the same as the local asymptotic behavior of the contour integral $I^{\varepsilon}_{1,1}$ when $x>\nicefrac{{1}}{{2}}$. [^5] This means if one is $O(N^{-1})$, then so is the other. Constant factors may be differing. The same is true for $I^{\varepsilon}_{2,1}$ and $I^{\varepsilon}_{1,2}$. Instead, to examine all possible cases of model behavior we will consider the following two cases.
  1. (a)
    $10<\beta<\alpha<$ 1
  2. (b)
    $10<\alpha<\beta<$ 1
This is because The behavior of $I_{2,2}^{\alpha}$ not only depends on the macroscopic regions of the model, but, if we are in case (b) above, it also depends on the region of the model in other ways. Consider the regions of the model where $(x,y)$ satisfies one of the following inequalities:
@@
    $\psi^{\prime}_{\alpha,2}(-\beta^{2};x,y)>0$
@@
    $\psi^{\prime}_{\alpha,2}(-\beta^{-2};x,y)<0$
Since $\psi^{\prime}_{\alpha,2}(z)$ is linear in $x$ and $y$, the inequalities just denote trapezoidal regions of the model. These regions are empty if $\beta<\alpha$. The exact size and shape of these regions depends on the exact values of $\alpha$ and $\beta$. Figure 8 gives examples of these regions for different values of $\alpha$ and $\beta$.
@@
We will denote the asymptotic region where either (i) or (ii) is satisfied as the strong-coupling region of the model, and denote it by $\mathfrak{C}$. We state the following asymptotic results regarding the behavior of $I^{\alpha}_{2,2}$,
@@
    $|I^{\alpha}_{2,2}(4m^{\prime},2\xi^{\prime}+j;4m,2\xi+i)|\leq cN^{-1}$
@@
    $|I^{\alpha}_{2,2}(4m^{\prime},2\xi^{\prime}+j;4m,2\xi+i)|\leq ce^{-c_{1}N}$
@@
    $|I^{\alpha}_{2,2}(4m^{\prime},2\xi^{\prime}+j;4m,2\xi+i)|\leq cN^{-1}$
@@
The behavior of the second double contour integral, $I^{\alpha}_{2,1}$, is less complicated as it does not depend on the macroscopic region of the model. We state the following proposition describing its behavior,
@@
$|I^{\alpha}_{2,1}(4m^{\prime},2\xi^{\prime}+j;4m,2\xi+i)|\leq ce^{-c_{1}N}$
@@
Lastly, we should mention how these results combine with the local asymptotic behavior of the typical two-periodic kernel. The local asymptotics for the two-periodic Aztec diamond were initially stated in [7, Theorem 2.6]. Let $\mathbb{K}_{FP}^{\varepsilon}(\bar{p}_{1};\bar{p}_{2})$ denote the full plane correlation kernel for the infinite square lattice with the same two-periodic weighting as the two-periodic Aztec diamond described in Section 2.2. The results from [7, Theorem 2.6] and also [11, Theorem 2.10] state [^6] The original results in [7] just focus along a diagonal of the model, the later results [11] look at the entire smooth region. that if $(x,y)\in\mathfrak{S}$ then,
(28) $\mathbb{K}^{\varepsilon}_{TP,N}(4m^{\prime},2\xi^{\prime}+j;4m,2\xi+i)=\mathbb{K}_{FP}^{\varepsilon}(x_{1},y_{1};x_{2},y_{2})+O(e^{-cN})$
However, in the case of the split two-periodic Asymptotic we see a difference in the sub leading order term. In particular, we wist to highlight that if $(x,y)\in\mathfrak{S}\cap\mathfrak{C}$ then
(29) $\mathbb{K}^{\varepsilon}_{TP,N}(4m^{\prime},2\xi^{\prime}+j;4m,2\xi+i)=\mathbb{K}_{FP}^{\varepsilon}(x_{1},y_{1};x_{2},y_{2})+O(N^{-1})$
In the region $\mathfrak{S}\cap\mathfrak{C}$, not only do the additional integrals contribute to the sub-leading order term, but they change the decay of it significantly.

### v5

@@ 2.1. An Integral Representation of the Correlation Kernel
Remark 2.1.
This is not the complete version of the kernel as it only accounts for vertices of the form $(4m,2\xi+i)$. In particular, there is a restriction on the horizontal coordinate being divisible by four. One can extend this kernel to account for all vertices by appending by certain matrix products that depend on the coordinates. For example, in the case where $m^{\prime}\leq\nicefrac{{N}}{{2}}$ we can extend the kernel by writing,
(14) $\Big[\mathbb{K}_{N}(4m^{\prime}-k,2\xi^{\prime}+j;4m+l,2\xi+i)\Big]_{i,j=0}^{1}\\ =-\frac{\mathbb{I}_{m>m^{\prime}}}{2\pi\textrm{i}}\oint_{\gamma_{0,1}}\frac{\text{d}z}{z}z^{\xi^{\prime}-\xi}T_{k,\alpha}(z)\Phi_{\alpha}(z)^{\frac{N}{2}-m^{\prime}}\Phi_{\varepsilon}(z)^{m-\frac{N}{2}}S_{l,\varepsilon}(z)\\ +\frac{1}{(2\pi\textrm{i})^{2}}\oint_{\gamma_{1}}\text{d}w\oint_{\gamma_{0,1}}\frac{\text{d}z}{z(z-w)}\frac{w^{\xi^{\prime}+N}(z-1)^{N}}{z^{\xi+N}(w-1)^{N}}r_{\alpha,1}(w)^{\frac{N}{2}-m^{\prime}}T_{k,\alpha}(w)F_{\alpha,1}(w)\Phi_{\varepsilon}(z)^{m-\frac{N}{2}}S_{l,\varepsilon}(z)\\ +\frac{1}{(2\pi\textrm{i})^{2}}\oint_{\gamma_{1}}\text{d}w\oint_{\gamma_{0,1}}\frac{\text{d}z}{z(z-w)}\frac{w^{\xi^{\prime}+N}(z-1)^{N}}{z^{\xi+N}(w-1)^{N}}r_{\alpha,2}(w)^{\frac{N}{2}-m^{\prime}}T_{k,\alpha}(w)\frac{2F_{\alpha,2}(w)F_{\beta,1}(w)F_{\alpha,1}(w)}{1+2g_{\alpha,\beta}(w)}\\ \times\Phi_{\varepsilon}(z)^{m-\frac{N}{2}}S_{l,\varepsilon}(z)$
where
$S_{0,\varepsilon}(z)=\mathbb{I},\;S_{1,\varepsilon}(z)=\phi_{\varepsilon,1}(z),\;S_{2,\varepsilon}(z)=\phi_{\varepsilon,1}(z)\phi_{\varepsilon,2}(z),\;S_{3,\varepsilon}(z)=\phi_{\varepsilon,1}(z)\phi_{\varepsilon,2}(z)\phi_{3}(z)$
and
$T_{0,\varepsilon}(z)=\mathbb{I},\;T_{1,\varepsilon}(z)=\phi_{4}(z),\;T_{2,\varepsilon}(z)=\phi_{3}(z)\phi_{4}(z),\;T_{3,\varepsilon}(z)=\phi_{\varepsilon,2}(z)\phi_{3}(z)\phi_{4}(z).$
Now the kernel can be computed for any vertex. In this work, we will only use the limited version of the kernel presented in Theorem 2.1 instead of this longer version. This is because the limited kernel has enough detail for understanding the asymptotic results of the model, without becoming overly excessive in notation. See [DK21, Remark 2.6] for a similar conversation regarding the kernel of the two-periodic Aztec diamond.
@@ 2.2. Comparison to the Two-Periodic Aztec Diamond
Before we analyze the behavior of the model, we first want to comment on how the form of the correlation kernel relates to the correlation kernel of the typical two-periodic Aztec diamond. From the definition of the model, we can recover the two-periodic Aztec diamond (with two-periodic parameter $\varepsilon$) from the split two-periodic Aztec diamond by setting $\alpha=\beta=\varepsilon$. The correlation kernel of the two-periodic Aztec diamond, as formulated in [BD19], can be stated using the notation presented in this paper as
Theorem ([BD19], Theorem 5.2).
Let $N\leq\xi,\xi^{\prime}\leq-1$ and $0<m<N$. The two-periodic Aztec diamond of size $2N$ has correlation kernel given by
(15) $\Big[\mathbb{K}_{TP,N}^{\varepsilon}(4m^{\prime},2\xi^{\prime}+j;4m,2\xi+i)\Big]_{i,j=0}^{1}=-\frac{\mathbb{I}_{m>m^{\prime}}}{2\pi\emph{i}}\oint_{\gamma_{0,1}}\frac{\text{d}z}{z}z^{\xi^{\prime}-\xi}\Phi_{\varepsilon}(z)^{m-m^{\prime}}\\ +\frac{1}{(2\pi\emph{i})^{2}}\sum_{k=1,2}\oint_{\gamma_{1}}\text{d}w\oint_{\gamma_{0,1}}\frac{\text{d}z}{z(z-w)}\frac{w^{\xi^{\prime}+N}(z-1)^{N}}{z^{\xi+N}(w-1)^{N}}r_{\varepsilon,1}(w)^{\frac{N}{2}-m^{\prime}}r_{\varepsilon,k}(z)^{m-\frac{N}{2}}F_{\varepsilon,1}(w)F_{\varepsilon,k}(z).$
Remark 2.2.
Once again, we have written the kernel only for coordinates of the form $(4m,2\xi+i)$. The kernel above can be extended to all coordinates via the method outlined in Remark 2.1. The matrix notation above is the same convention as the matrix notation in Theorem 2.1.
@@
(16) $\Big[\mathbb{K}_{N}(4m^{\prime},2\xi^{\prime}+j;4m,2\xi+i)\Big]_{i,j=0}^{1}=\Big[\mathbb{K}_{TP,N}^{\varepsilon}(4m^{\prime},2\xi^{\prime}+j;4m,2\xi+i)\Big]_{i,j=0}^{1}\\ +\frac{1}{(2\pi\textrm{i})^{2}}\sum_{k=1,2}\oint_{\gamma_{1}}\text{d}w\oint_{\gamma_{0,1}}\frac{\text{d}z}{z(z-w)}\frac{w^{\xi^{\prime}+N}(z-1)^{N}}{z^{\xi+N}(w-1)^{N}}r_{\varepsilon,l}(w)^{\frac{N}{2}-m^{\prime}}r_{\varepsilon,k}(z)^{m-\frac{N}{2}}\\ \times\frac{2}{1+2g_{\alpha,\beta}(w)}F_{\varepsilon,l}(w)F_{\bar{\varepsilon},1}(w)F_{\varepsilon,\bar{l}}(w)F_{\varepsilon,k}(z)\\$
@@
[…]
@@ 2.3. Classification of the Macroscopic Regions
Before we compute the explicit local asymptotics of the model, we would like to introduce the saddle functions associated with the kernel of the model. In doing so, we will consider the behavior of the model around some asymptotic coordinate. Let $(x,y)\in(0,\nicefrac{{1}}{{2}})\cup(\nicefrac{{1}}{{2}},1)\times(-1,0)$, where we omit the case where the asymptotic coordinate is exactly on the interface of the model. Let $(m,\xi)=(xN,yN)$, then we say the vertex $(4m,2\xi+i)$ has asymptotic coordinate (x,y)(x,y). [^3] Calling $(x,y)$ the asymptotic coordinate means that the asymptotic picture fits inside the square $0,1]\times[-1,0]$, which is the typical asymptotic view of the Aztec diamond. Now we define the functions,
@@
The saddle functions $\psi_{\varepsilon,k}(z)$ are, up to a scalar and change of parameters, the saddle functions defined and analyzed in [DK21].
In other words, most of the analytic work we need in order to prepare our integrals for saddle point methods was accomplished in [DK21]. We state the saddle functions in [DK21, p. 15] for comparison.
@@
$\lambda_{1,2}(z)=\frac{\left((\alpha_{DK}+\beta_{DK})z\pm\sqrt{z(z+\alpha_{DK}^{2})(z+\beta_{DK}^{2})}\right)^{2}}{z(z-1)^{2}}$
and $1<\xi_{1},\xi_{2}<1$ are parameters that define the asymptotic coordinate, albeit different from our $x$ and $y$. Note that the above $\alpha_{DK}$ and $\beta_{DK}$ are from the convention in [DK21] and not the same $\alpha$ and $\beta$ presented in this text. Letting $\alpha_{DK}=\varepsilon$ and $\beta_{DK}=\varepsilon^{-1}$ it is easy to check that,
@@
There are four saddle points, counting multiplicity, of the function $\Psi(z)$ as stated in [DK21]. The location of these saddle points on the Riemann surface classify the macroscopic region of the model,
@@
    If $\Psi$ has two simple saddle points on either copy of the positive real axis, then $(x,y)\in\mathfrak{F}$, the frozen region of the model.
@@
    If $\Psi$ has two complex saddle points (on either sheet of the Riemann surface), then $(x,y)\in\mathfrak{R}$, the rough region of the model.
@@
    If all four saddle points of $\Psi$ are simple and located on either copy of the negative real axis away from the branch cuts, then $(x,y)\in\mathfrak{S}$, the smooth region of the model.
This is the same definition of regions presented for the typical two-periodic model in [DK21, Definition 2.9]. Moreover, Duits and Kuijlaars go on to define the phase transitions as tuples, $(x,y)$, where the saddle points of the function coalesce. We do not state this full definition here, but the same applies as long as $x\neq\nicefrac{{1}}{{2}}$. The saddle function, $\Psi(z)$, can also be used to determine the boundaries between the macroscopic regions. Any point $(x,y)$ where saddle points coalesce is on some boundary between the macroscopic regions. Additionally, due to the piece wise nature of the saddle functions, we must check for transitions along the line $x=\nicefrac{{1}}{{2}}$. Figure 8 depicts an example of the macroscopic boundaries for a split two-periodic model.
@@
For either half of the model, the definition and boundaries between macroscopic regions agree with the typically two-periodic case. [^5] See [DK21] for a similar classification in the two periodic case. In the section below we will show how this follows from the local asymptotics of the correlation kernel.
@@ 2.4. Asymptotic Behavior
Our goal in this section is to summarize the local asymptotics of the model, so we will assume the two coordinates $(4m^{\prime},2\xi^{\prime}+j)$ and $(4m,2\xi+i)$ are on the same side of the interface and that they are asymptotically close. We define this precisely. Start by writing,
@@
where we assume $x\neq\nicefrac{{1}}{{2}}$, and $x_{1}$ and $x_{2}$ are small enough in magnitude so that the coordinates $(4m,2\xi+i)$ and $(4m^{\prime},2\xi^{\prime}+j)$ lie on the same side of the interface. In general, we will assume that $x_{1},\,x_{2},\,y_{1},$ and $y_{2}$ are $o(N)$. Under these considerations, we rewrite the correlation kernel in terms of the saddle functions,
(21) $\Big[\mathbb{K}_{N}(4m^{\prime},2\xi^{\prime}+j;4m,2\xi+i)\Big]_{i,j=0}^{1}=\Big[\mathbb{K}_{TP,N}^{\varepsilon}(4m^{\prime},2\xi^{\prime}+j;4m,2\xi+i)\Big]_{i,j=0}^{1}\\ +\frac{1}{(2\pi\textrm{i})^{2}}\sum_{k=1,2}\oint_{\gamma_{1}}\text{d}w\oint_{\gamma_{0,1}}\frac{\text{d}z}{z(z-w)}\frac{2}{1+2g_{\alpha,\beta}(w)}F_{\varepsilon,l}(w)F_{\bar{\varepsilon},1}(w)F_{\varepsilon,\bar{l}}(w)F_{\varepsilon,k}(z)\\ \times\exp\left[N\left(\psi_{l}(w;x,y)-\psi_{k}(z;x,y)\right)+\varphi_{\varepsilon,l}(w;x_{2},y_{2})-\varphi_{\varepsilon,k}(z;x_{1},y_{1})\right]$
@@
For brevity we let
(23) $\Big[I^{\varepsilon}_{l,k}(4m^{\prime},2\xi^{\prime}+j;4m,2\xi+i)\Big]_{i,j=0}^{1}\\ =\frac{1}{(2\pi\textrm{i})^{2}}\oint_{\gamma_{1}}\text{d}w\oint_{\gamma_{0,1}}\frac{\text{d}z}{z(z-w)}\frac{2}{1+2g_{\alpha,\beta}(w)}F_{\varepsilon,l}(w)F_{\bar{\varepsilon},1}(w)F_{\varepsilon,\bar{l}}(w)F_{\varepsilon,k}(z)\\ \times\exp\left[N\left(\psi_{l}(w;x,y)-\psi_{k}(z;x,y)\right)+\varphi_{\varepsilon,l}(w;x_{2},y_{2})-\varphi_{\varepsilon,k}(z;x_{1},y_{1})\right]$
for $k=1,2$. So we may write the kernel, compactly, as
@@
Ultimately, our goal is to show that the contributions from the contour integrals $I^{\alpha}_{2,1}$ and $I^{\alpha}_{2,2}$ tend to 0 as $N\to\infty$. This will justify the definition of the macroscopic boundary given in Definition 2.1. The order of decay depends on the region of the model of the asymptotic coordinate. We will find that, in certain regions, these additional terms affect the second order term of the local asymptotics of $\mathbb{K}_{N}$.
@@
which means there is a symmetry between the saddle functions when the lattice is flipped along the interface. This symmetry is not enough for equality between the contour integrals, but it is enough, along with the proofs detailed in Section 4, to conclude that the asymptotic behavior is the same. The difference in matrices still present in the integral only change a constant factor, not the rate of decay. If we let $m$, $m^{\prime}$, $\xi$, $\xi^{\prime}$ be defined by equations (19) and (20) and $m,\,m^{\prime}<N/2$, then $I^{\varepsilon}_{2,2}(4m^{\prime},2\xi^{\prime}+j;4m,2\xi+i)$ has the same asymptotic behavior as $I^{\varepsilon}_{1,2}(4(N-m^{\prime}),2\xi^{\prime}+j;4(N-m),2\xi+i)$. Additionally, $I^{\varepsilon}_{2,1}(4m^{\prime},2\xi^{\prime}+j;4m,2\xi+i)$ has the same asymptotic behavior as $I^{\varepsilon}_{1,1}(4(N-m^{\prime}),2\xi^{\prime}+j;4(N-m),2\xi+i)$. Thus below we only state asymptotic results for $I_{2,2}^{\alpha}$ and $I_{2,1}^{\alpha}$. The asymptotics on the $\beta$-side of the interface can be proven in an identical manner.
Remark 2.4.
There is something interesting going on with regard to the symmetry of the model. There is not a discrete symmetry, like in the two-periodic model. One could try to switch parameters, flip the lattice, and then gauge transform but the interface becomes a problem and there exists no gauge transform. The is, however, a symmetry of the limit shape because the additional integrals (compared to the two-periodic model) behave the same, up to a scalar, asymptotically.
To examine all possible cases of model behavior we will consider the following two cases: (a) $0<\beta<\alpha<1$, and (b) $0<\alpha<\beta<1$. This is because the asymptotic behavior of $I_{2,2}^{\alpha}$ not only depends on the macroscopic regions of the model, but, if we are in case (b) above, it also depends on the region of the model in other ways. We restrict to the case of $x<\nicefrac{{1}}{{2}}$ and case (b) above. We can then consider the regions of the model where $(x,y)$ satisfies one of the following inequalities:
@@
    $\psi^{\prime}_{\alpha,1}(-\beta^{2};x,y)>0$
@@
    $\psi^{\prime}_{\alpha,1}(-\beta^{-2};x,y)<0$
Since $\psi^{\prime}_{\alpha,1}(z)$ is linear in $x$ and $y$, the inequalities just denote trapezoidal regions of the model. As we will show later in Lemma 4.2, $\operatorname{Re}\psi_{\alpha,1}(z)$ has exactly one maximum on the interval $\alpha^{-2},-\alpha^{2}]$, thus we conclude that no $(x,y)$ satisfies both inequalities. In the case of $\beta<\alpha$ we define no such region. The exact size and shape of these regions depends on the exact values of $\alpha$ and $\beta$. Figure 9 gives examples of these regions for different values of $\alpha$ and $\beta$.
@@
We will denote the asymptotic region where either (i) or (ii) is satisfied as the strong-coupling region of the model, and denote it by $\mathfrak{C}$. As we will see below, the coupling between the two weightings of the model is strongest in this region. We state the following asymptotic results regarding the behavior of $I^{\alpha}_{2,2}$,
@@
    $|I^{\alpha}_{2,2}(4m^{\prime},2\xi^{\prime}+j;4m,2\xi+i)|=cN^{-1}+o(N^{-1})$
@@
    $|I^{\alpha}_{2,2}(4m^{\prime},2\xi^{\prime}+j;4m,2\xi+i)|=ce^{-c_{1}N}+o(e^{-c_{1}N})$
@@
    $|I^{\alpha}_{2,2}(4m^{\prime},2\xi^{\prime}+j;4m,2\xi+i)|=cN^{-1}+o(N^{-1})$
@@
The exact constants can be extracted from the proof in Section 4. The behavior of the second double contour integral, $I^{\alpha}_{2,1}$, is less complicated as it does not depend on the macroscopic region of the model. We state the following proposition describing its behavior,
@@
$|I^{\alpha}_{2,1}(4m^{\prime},2\xi^{\prime}+j;4m,2\xi+i)|=ce^{-c_{1}N}+o(e^{-c_{1}N})$
@@
Moreover, the inequalities presented in the above propositions are optimal. Lastly, we should mention how these results combine with the local asymptotic behavior of the typical two-periodic kernel. The local asymptotics for the two-periodic Aztec diamond were initially stated in [CJ16, Theorem 2.6]. Let $\mathbb{K}_{FP}^{\varepsilon}(\bar{p}_{1};\bar{p}_{2})$ denote the full plane correlation kernel for the infinite square lattice with the same two-periodic weighting as the two-periodic Aztec diamond described in Section 2.2. The results from [CJ16, Theorem 2.6] and also [DK21, Theorem 2.10] state [^6] The original results in [CJ16] just focus along a diagonal of the model, the later results [DK21] look at the entire smooth region. that if $(x,y)\in\mathfrak{S}$ then,
(26) $\mathbb{K}^{\varepsilon}_{TP,N}(4m^{\prime},2\xi^{\prime}+j;4m,2\xi+i)=\mathbb{K}_{FP}^{\varepsilon}(x_{1},y_{1};x_{2},y_{2})+ce^{-c_{1}N}+O(e^{-c_{1}N}),$
where $c$ is a non-zero constant. However, in the case of the split two-periodic Asymptotic we see a difference in the sub leading order term. In particular, we wist to highlight that if $(x,y)\in\mathfrak{S}\cap\mathfrak{C}$ then
(27) $\mathbb{K}^{\varepsilon}_{TP,N}(4m^{\prime},2\xi^{\prime}+j;4m,2\xi+i)=\mathbb{K}_{FP}^{\varepsilon}(x_{1},y_{1};x_{2},y_{2})+cN^{-1}+O(N^{-1}),$
where $c$ is again a non-zero constant. In the region $\mathfrak{S}\cap\mathfrak{C}$, not only do the additional integrals contribute to the sub-leading order term, but they change the decay of it significantly. Note that this does not happen in $\mathfrak{R}\cap\mathfrak{C}$, because in the rough region the two-periodic kernel converges to the related full plane kernel with an error of $N^{-1/2}$.